Peavey Enterprises purchased a depreciable asset for $32,000 on April 1, Year 1. The asset will be depreciated using the straight-line method over its four-year useful life. Assuming the asset's salvage value is $4,000, what will be the amount of accumulated depreciation on this asset on December 31, Year 3?

Multiple Choice

$19,250

$5,833

$5,833

$7,000

$23,333
Answer #1

Solution:

Annual Depreciation = (cost - salvage) / 4 = ($32000 - $4000) / 4 = $7,000

Depreciation for Year 1 (for 9 months) = $7000*9/12 = $5,250

Accumulated Depreciation on December 31, Year 3 = $5250 + $7000 + $7000 = $19,250

Hence first option is correct.